OpenStudy (anonymous):

change y=3x^2 +1 into polar form

OpenStudy (anonymous):

what do x,y equal in polar cord? \[x=r\cos\theta \text{ and } y = ?\]

OpenStudy (anonymous):

\[y=r \sin \theta\]

OpenStudy (anonymous):

@Redcan

OpenStudy (anonymous):

ok, now sub the new equations into the original

OpenStudy (anonymous):

\[rsin \theta =3\cos \theta ^2+1\]

OpenStudy (anonymous):

Yep, \[r\sin\theta=3(r\cos\theta)^2+1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

now you can simplify if you want... recall \[sin^2\theta + \cos^2\theta = 1\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so replace the cos^2 with 1

OpenStudy (anonymous):

Um, nope. we would need a sin^2 as well to do that.